if a car has a mass of 200 kg moving in a circular path with velocity 10m/s and the centripetal force acting on it is 2000 newtons whats the radius and acceleration
1
Expert's answer
2020-05-21T14:00:01-0400

"F=ma \rightsquigarrow=a=F/m=2000/200=10\,m/s"

"a=\frac{v^2}{R} \rightsquigarrow R=v^2/a=10^2/10=10\,m"